A rectangle is drawn so that the width is 4 feet shorter than the length. The area of the rectangle is 60 square feet. Find the length of the rectangle.
Answer by Boreal(15235)

width is x
length is x+4
product is area of x^2+4x, and that equals 60
so x^2+4x-60=0
(x+10)(x-6)=0
x=6 feet, only positive root
x+4=10 feet length ANSWER